Explain why the Hall effect occurs when a wire with a square cross-section of side length w carrying charges moving with a drift velocity vd is placed in a uniform magnetic field B orientated perpendicular to the side of the wire. Give an expression for the magnitude of the Hall voltage produced in this case


1
Expert's answer
2021-05-17T19:00:02-0400

Gives

Drift velocity J=neVd

we know that


hall constant RH"=\frac{E_H}{J_xB_z}\rightarrow(1)"

"R_H=\frac{1}{nq}\rightarrow(2)"

Then we know that

eqution (1)and(2) we can write as

"R_H=\frac{1}{nq}=\frac{E_H}{J_xB_z}"

Wire is square type

Hall electric field "E_H=\frac{V_H}{w}" "\rightarrow(3)"

"E_H=R_HJ_xB_z""\rightarrow(4)"

Current density "J_x=\frac{I_x}{A}\rightarrow(5)"

eqution(3)(4)and(5) we can write as

Then hall voltage

"V_H=E_Hw"

w=sample weight